A solid body starts rotating about a stationary axis with an angular acceleration  a = (2 x 10 ^ -2)t rad/second squared., here t is in seconds. How soon after the beginning of rotation will the total acceleration vector of an arbitrary point of the body form an angle theta = 60 degrees with its velocity vector?

Grade:12th Pass

2 Answers

Chetan Mandayam Nayakar
11 years ago

tan(600)=(v2/r)/(ra)

If you want more explanation, please contact me.

RAJEEV CHOUDHARY
11 years ago

T^3 = 50(3^0.5)